# Benchmarking XAI Explanations with Human-Aligned Evaluations

### v1

@@ A.3 Dataset Annotation process
Each participant’s task was to annotate 147 explanations. For each image, participants had to explain what led them to classify the displayed image as the model’s prediction. Participants responded openly using a text form. Similarly, they were asked to describe the elements of the image that helped them make the decision to classify the image as the model did. These two questions (Q0.1 and Q0.2) are used to establish a baseline for interpreting what makes an image recognizable as a specific object or class, and what are the salient features of the images that would explain this choice.
Once they had answered these two questions, they moved on to the next screen, where they had to answer a series of 6 questions (see Figure 9 showing a screenshot of the interface with the original image slightly perturbed). The questions concerned different sections of the screen. The first four questions (Q1 to Q4) concerned Section 1, which showed the original image on the left and the explanation on the right. These first four questions enabled participants to assess the levels of reliability, complexity and objectivity of the explanation. The fifth question (Q5) concerned Section 2, showing the slightly disturbed original image and the corresponding explanation. The sixth question (Q6) concerned Section 3, showing a more disturbed image and the corresponding explanation. These last two questions were intended to assess the robustness of the XAI technique. For each question, participants had to answer with a 5-point Likert scale.
All questions were formulated in collaboration with psychologists to enhance the quality of human feedback. Our aim was to avoid influencing responses and to eliminate any ambiguities that could lead to inaccurate answers. Human decision-making, especially when it involves assessing the quality of explanations, is complex. To address this, we provided training on deep learning and various XAI techniques, ensuring the content was clearly understandable by the annotators. After the initial training, the annotators answered the questions, and we held weekly meetings to clarify any confusion they encountered
@@
Appendix B PASTA-dataset: additional experiments
B.1 Comparison with existing metrics
Evaluating the quality of an explanation typically involves estimating different and potentially orthogonal aspects of it. In addition to the perceptual quality addressed in this work, others can be numerically simulated by having access to model weights. In this additional analysis, we consider some of those aspects and measure how much they correlate with human scores. The results cover only image-level attribution methods (see Table 6), as CBMs do not support such kinds of input-level manipulations.
@@ B.1 Comparison with existing metrics
How much does the explanation describe the true behavior of the model? A number of different ways to compute faithfulness exist, but they all broadly fit the same framework of measuring how much model predictions change in response to input perturbations (Bhatt et al., 2021; Alvarez-Melis & Jaakkola, 2018a; Yeh et al., 2019; Rieger & Hansen, 2020; Arya et al., 2019; Nguyen & Martínez, 2020; Bach et al., 2015; Samek et al., 2016; Montavon et al., 2018; Ancona et al., 2017; Dasgupta et al., 2022). Intuitively, an explanation is faithful if perturbing regions deemed irrelevant by the explanation bring little to no change in model output, whereas perturbing regions deemed relevant bring a considerable change. In this analysis, we resort to the evaluation protocol outlined in Azzolin et al. (2024), which generalized a number of common faithfulness metrics into a common mold [^1] They focus on faithfulness for graph explanations, but the evaluation protocol is aligned with that of images.. Specifically, faithfulness is estimated as the harmonic mean of sufficiency and necessity, which account for the degree of prediction changes after perturbing irrelevant or relevant portions of the input, respectively. There is no unique way of applying perturbations to an image, and different techniques are oftentimes reported to give different interpretations (Hase et al., 2024; Rong et al., 2022). To avoid this confounding effect, we report the results for three different baseline perturbations, namely uniform and Gaussian noise, and black patches, along with a more advanced information-theoretic strategy named ROAD (Rong et al., 2022). Since explanations are oftentimes in the form of soft relevance scores over the entire input, a threshold is needed to tell apart relevant from irrelevant image regions. To avoid relying upon this hard-to-define hyperparameter, we aggregate the scores across multiple thresholds keeping only the best value. Therefore, for each threshold value, pixels are sorted based on their relevance [^2] For sufficiency, pixels are sorted in ascending order. For necessity, in descending order. and progressively perturbed until reaching the fixed threshold value, while leaving the others unchanged. Then, we measure the absolute difference in class-predicted confidence between clean and perturbed images, normalizing the results to ensure that both are the higher the better. The area under the curve is finally computed, indicating how fast the model changes its predictions with respect to the amount of perturbed pixels.
@@
Robustness roughly refers to how stable the explanation is to small input perturbations. Different ways to estimate it exist (Alvarez-Melis & Jaakkola, 2018b; Montavon et al., 2018; Yeh et al., 2019; Dasgupta et al., 2022; Agarwal et al., 2022a). In our analysis, we focused on MaxSensitivity (Yeh et al., 2019), which applies random input perturbations to the entire image and measures the pixel-wise difference between the original explanation, and the one obtained on the perturbed sample. Again, different perturbation techniques can be applied, and we resort to the two simple baselines, namely uniform and Gaussian noise. No normalization is applied, therefore the values are the higher the worse. More advanced techniques like ROAD (Rong et al., 2022) cannot be applied in this context, since the perturbation is applied uniformly over the entire image. In Table 7, we report the correlation between MaxSensitivity and human scores, outlining a non-significant correlation with the metric and some questions. Surprisingly, the most correlated questions are Q1-4, which are not requesting humans to assess the stability of the explanation, something instead partially addressed by Q5 and Q6. However, the correlation is very weak anyway, questioning any further claims.
@@
As humans have an implicit tendency to favor simple alternatives when facing a comparison between different hypotheses, providing simple and compact explanations is vital for human-machine synergy (Cowan, 2001). Alternative methods for estimating the complexity of an explanation are available, from simple above-threshold counting to more advanced information-theoretic techniques (Chalasani et al., 2020; Bhatt et al., 2021; Nguyen & Martínez, 2020). To test whether those metrics are correlated to human scores, we report in Table 8 the correlation between human votes and Sparseness (Chalasani et al., 2020), which estimates explanation complexity as the Gini Index (Hurley & Rickard, 2009) of the absolute values of the image attribution. The result is a metric value in the range $[0,1]$, where higher values indicate more sparseness.
@@
Q1 -0.17 (1e-10) -0.15 (1e-8)
Q2 -0.17 (1e-10) -0.15 (1e-8)
Q3 -0.15 (1e-9) -0.13 (1e-7)
Q4 -0.15 (1e-7) -0.13 (1e-6)
Q5 -0.21 (1e-15) -0.22 (1e-17)
Q6 0.02 (0.45) 0.05 (0.09)
We used the Quantus library (Hedström et al., 2023) for implementing the previous metrics, and we present the raw metric values in Table 9, aggregated by explainer and model. Overall, none of the above metrics exhibit a significant correlation with user scores.
@@
GradCAM resnet50 0.08 $\pm$ 0.16 0.79 $\pm$ 0.45 0.68 $\pm$ 0.11
GradCAM vitB 0.02 $\pm$ 0.04 1.11 $\pm$ 0.96 0.41 $\pm$ 0.19
AblationCAM vitB 0.02 $\pm$ 0.02 1.83 $\pm$ 1.68 0.77 $\pm$ 0.23
AblationCAM CLIP-zero-shot 0.01 $\pm$ 0.01 1.46 $\pm$ 0.44 0.83 $\pm$ 0.15
EigenCAM resnet50 0.11 $\pm$ 0.19 0.94 $\pm$ 0.46 0.80 $\pm$ 0.05
EigenCAM vitB 0.02 $\pm$ 0.03 0.98 $\pm$ 0.35 0.59 $\pm$ 0.06
EigenCAM CLIP-zero-shot 0.02 $\pm$ 0.02 0.76 $\pm$ 0.37 0.54 $\pm$ 0.08
EigenGradCAM resnet50 0.11 $\pm$ 0.18 0.94 $\pm$ 0.70 0.78 $\pm$ 0.10
EigenGradCAM vitB 0.02 $\pm$ 0.03 2.10 $\pm$ 1.38 0.88 $\pm$ 0.11
EigenGradCAM CLIP-zero-shot 0.03 $\pm$ 0.02 1.55 $\pm$ 1.27 0.73 $\pm$ 0.17
FullGrad resnet50 0.05 $\pm$ 0.14 0.43 $\pm$ 0.06 0.43 $\pm$ 0.07
FullGrad vitB 0.02 $\pm$ 0.02 1.30 $\pm$ 0.46 0.39 $\pm$ 0.06
GradCAM CLIP-zero-shot 0.03 $\pm$ 0.03 1.29 $\pm$ 0.65 0.60 $\pm$ 0.16
GradCAMPlusPlus resnet50 0.08 $\pm$ 0.13 0.62 $\pm$ 0.20 0.63 $\pm$ 0.09
GradCAMPlusPlus vitB 0.02 $\pm$ 0.03 2.45 $\pm$ 1.90 0.69 $\pm$ 0.24
GradCAMPlusPlus CLIP-zero-shot 0.01 $\pm$ 0.01 1.53 $\pm$ 0.97 0.65 $\pm$ 0.23
GradCAMElementWise resnet50 0.07 $\pm$ 0.16 0.70 $\pm$ 0.34 0.58 $\pm$ 0.08
GradCAMElementWise vitB 0.02 $\pm$ 0.03 1.17 $\pm$ 0.37 0.61 $\pm$ 0.11
GradCAMElementWise CLIP-zero-shot 0.03 $\pm$ 0.03 0.73 $\pm$ 0.16 0.38 $\pm$ 0.07
HiResCAM resnet50 0.12 $\pm$ 0.19 0.86 $\pm$ 0.32 0.65 $\pm$ 0.11
HiResCAM vitB 0.02 $\pm$ 0.02 1.71 $\pm$ 0.83 0.70 $\pm$ 0.22
HiResCAM CLIP-zero-shot 0.02 $\pm$ 0.03 1.45 $\pm$ 0.41 0.68 $\pm$ 0.13
LIME resnet50 0.13 $\pm$ 0.20 0.50 $\pm$ 0.07 0.12 $\pm$ 0.04
ScoreCAM resnet50 0.07 $\pm$ 0.17 0.85 $\pm$ 0.43 0.56 $\pm$ 0.09
ScoreCAM vitB 0.03 $\pm$ 0.04 1.24 $\pm$ 0.80 0.46 $\pm$ 0.18
XGradCAM resnet50 0.14 $\pm$ 0.19 1.00 $\pm$ 0.41 0.71 $\pm$ 0.10
XGradCAM vitB 0.02 $\pm$ 0.02 1.32 $\pm$ 0.16 0.61 $\pm$ 0.06
XGradCAM CLIP-zero-shot 0.03 $\pm$ 0.03 1.31 $\pm$ 0.12 0.58 $\pm$ 0.04
DeepFeatureFactorization resnet50 0.07 $\pm$ 0.17 1.28 $\pm$ 0.54 0.34 $\pm$ 0.11
DeepFeatureFactorization vitB 0.02 $\pm$ 0.03 0.62 $\pm$ 0.21 0.28 $\pm$ 0.05
DeepFeatureFactorization CLIP-zero-shot 0.03 $\pm$ 0.05 0.73 $\pm$ 0.36 0.36 $\pm$ 0.09
@@
BCos resnet50-bcos 0.17 $\pm$ 0.25 0.85 $\pm$ 0.32 0.50 $\pm$ 0.09
LIME CLIP-zero-shot 0.03 $\pm$ 0.04 0.61 $\pm$ 0.10 0.16 $\pm$ 0.05
SHAP resnet50 0.12 $\pm$ 0.20 1.38 $\pm$ 0.55 0.50 $\pm$ 0.11
SHAP vitB 0.02 $\pm$ 0.03 1.05 $\pm$ 0.48 0.40 $\pm$ 0.10
SHAP CLIP-zero-shot 0.02 $\pm$ 0.03 1.31 $\pm$ 0.52 0.48 $\pm$ 0.11
AblationCAM resnet50 0.12 $\pm$ 0.22 0.95 $\pm$ 0.48 0.59 $\pm$ 0.12
B.2 Dataset analysis
To thoroughly analyze the dataset and evaluate potential biases, we conducted several tests. First, we experimented with various aggregation techniques, ultimately selecting the majority voting method as the most effective.

### v2

@@ A.2 Human evaluation protocol
Concretely, annotators are shown a succession of samples that consist in an image, a prediction, and an explanation. They are then asked to respond to a set of questions corresponding to the desiderata outlined above (indicated in italics):
  • Q1: Is the provided explanation consistent with how I would explain the predicted class? Trustworthiness
  • Q2: Overall the explanation provided for the model prediction can be trusted? Trustworthiness
  • Q3: Is the explanation easy to understand? Complexity
  • Q4: Can the explanation be understood by a large number of people, independently of their demographics (age, gender, country, etc.) and culture? Objectivity
A screenshot of the interface used by the annotators to answer is available in Figure 14. The first four questions (Q1 to Q4) concerned Section 1, which showed the original image on the left and the explanation on the right. These first four questions enabled participants to assess the levels of reliability, complexity and objectivity of the explanation. The fifth question (Q5) concerned Section 2, showing the slightly disturbed original image and the corresponding explanation. The sixth question (Q6) concerned Section 3, showing a more disturbed image and the corresponding explanation. These last two questions were intended to assess the robustness of the XAI technique. For each question, participants had to answer with a 5-point Likert scale. In question 5, we apply a weak data augmentation to the input image, which is designed to preserve the classifier’s prediction. This allows us to evaluate whether the explanation changes when the prediction remains constant. Measuring changes in explanations is challenging, as studied by (Fel et al. 2022b). To address this, we leverage human evaluators who can more effectively discern subtle changes in explanations. Q6 follow a similar approach but involve strong data augmentation. The goal here is to determine if the explanations remain consistent when the predictions change due to data augmentation or variations in the models. If the predictions change, the explanations should reflect these changes. For more details, refer to Figure 15.
@@
We outline the perturbation process below. To capture a diverse range of model responses, we applied 12 distinct types of perturbations, each with a tunable magnitude parameter to adjust perturbation intensity. The transformations include both standard torchvision operations (https://pytorch.org/vision/0.9/transforms.html) and custom-designed modifications:
  • Color Jitter (Brightness): Adjusts the brightness of the image. The magnitude lower the brightness.
  • Color Jitter (Contrast): Modifies image contrast. The magnitude lower the contrast.
  • Random Resized Crop: Performs a random crop and resize. The magnitude augments the scale of the crop.
  • Gaussian Blur: Blurs the image using a Gaussian filter. The magnitude augments the values of the standard deviation.
  • Random Perspective: Applies a perspective transformation. The magnitude augments the distortion scale.
  • Brightness Transform: Independently changes brightness levels. The magnitude lower the brightness.
  • Color Transform: Adjusts color balance. The magnitude augments the saturation.
  • Contrast Transform: Further modifies contrast. The magnitude augments the contrast.
  • Sharpness Transform: Changes image sharpness. The magnitude augments the sharpness factor.
  • Posterize Transform: Reduces color depth. The magnitude augments the number of bits to keep for each channel.
  • Solarize Transform: Inverts colors above a certain threshold. The magnitude augments the threshold.
  • Random Masking: Masks out random sections of the image by applying patches. The magnitude augments the number of patches.
Our aim was to avoid influencing responses and to eliminate any ambiguities that could lead to inaccurate answers. Human decision-making, especially when it involves assessing the quality of explanations, is complex. To address this, we provided training on deep learning and various XAI techniques, ensuring the content was clearly understandable by the annotators. After the initial training, the annotators answered the questions, and we held weekly meetings to clarify any confusion they encountered.
(a) A light perturbation is applied, resulting in no label change.
(b) A strong perturbation is applied, leading to a label change.
Figure 15: Samples corresponding to questions 5 and 6.
Each explanation is rated on a five points Likert scale, where one star indicates a total disagreement with the annotator’s reasoning, and five stars represent total agreement.
Appendix B B. PASTA-dataset: Additional analysis
B.1 Comparison with existing benchmark
We compiled a set of related works that perform human assessment in the context of XAI. Specifically, we noted key details such as the dataset size, the number of participants involved, the diversity of questions posed, and the overall scope of the study, where this information was available. A summary of these details is presented in Table 6.
The PASTA-dataset distinguishes itself in several key aspects. First, it evaluates a significantly higher number of XAI methods compared to existing datasets. This design emphasizes the diversity of techniques over the number of samples tested, offering a complementary approach to datasets that prioritize varied input data but evaluate fewer methods. Second, PASTA involves the lowest number of participants among the datasets listed, allowing for reduced variability due to potential outliers and better control over annotator behavior. However, this could introduce inherent biases tied to the limited participant pool. Finally, the PASTA-dataset provides a substantially larger number of samples and uniquely combines image-based explanations with concept-based explanations, making it the first dataset to address both modalities simultaneously.
Table 6: Overview of datasets and human evaluation frameworks for XAI methods. Name refers to the reference of the dataset used. Annotations refers to the type of labels used: Likert refers to Likert scale, Saliency refers to pseudo saliency maps, 2AFC refers to two alternative forced choices, Clictionary refers to the clicktionary game defined in (Dawoud et al. 2023), MCQ refers to multiple-choice question, and Binary refers to binary choises. $N_{Samples}$ refers to the total number of samples constituting the dataset. $N_{Part}$ refers to the number of participants involved during the labeling. Modality refers to the different modalities the dataset deals with: I refers to image, C to concepts, and T to text. $N_{Q}$ refers to the number of different questions asked to the annotators. $N_{XAI}$ refers to the number of XAI methods tested during the experiments, No indicates that the dataset only asked to label data with what they consider as ground truth explanation, without further comparison with XAI methods. $N_{Data}$ refers to the number of different samples (for example images) shown to annotators.
Name Annotations $N_{Samples}$ $N_{Part}$ Modality $N_{Q}$ $N_{XAI}$ $N_{Data}$
PASTA-dataset Likert 633,000 24 I + C 6 20 1000
Yang, Folke, and Shafto (2022) Saliency, 2AFC 356 46 I 2 1 89
Colin et al. (2022) Classification 1,960 241 I 1 6 NA
Dawoud et al. (2023) Clicktionary 3,836 76 I 1 3 102
Mohseni, Block, and Ragan (2021) Saliency 1,500 200 I + T 1 No 1,500
Herm et al. (2021) Likert NA 165 C 1 6 NA
Morrison et al. (2023) Clicktionary/QCM 450 50 I 1 3 39
Spreitzer, Haned, and van der Linden (2022) Likert/Binary 4,050 135 C 9 2 NA
Xuan et al. (2023) Likert/Binary 3,600 200 C 4 2 1,326
B.2 Comparison with existing metrics
@@
How much does the explanation describe the true behavior of the model? A number of different ways to compute faithfulness exist, but they all broadly fit the same framework of measuring how much model predictions change in response to input perturbations (Bhatt, Weller, and Moura 2021; Alvarez-Melis and Jaakkola 2018a; Yeh et al. 2019; Rieger and Hansen 2020; Arya et al. 2019; Nguyen and Mart´ınez 2020; Bach et al. 2015; Samek et al. 2016; Montavon, Samek, and Müller 2018; Ancona et al. 2017; Dasgupta, Frost, and Moshkovitz 2022). Intuitively, an explanation is faithful if perturbing regions deemed irrelevant by the explanation bring little to no change in model output, whereas perturbing regions deemed relevant bring a considerable change. In this analysis, we resort to the evaluation protocol outlined in Azzolin et al. (2025), which generalized a number of common faithfulness metrics into a common mold [^1] They focus on faithfulness for graph explanations, but the evaluation protocol is aligned with that of images.. Specifically, faithfulness is estimated as the harmonic mean of sufficiency ($\mathsf{Suf}$) and necessity ($\mathsf{Nec}$), which account for the degree of prediction changes after perturbing irrelevant or relevant portions of the input, respectively. Formally, given an input image $\bm{x}$} with associated explanation $\bm{e}$}, and a model to be explained $p_{\theta}(Y\mid{\bm{x}})$, sufficiency and necessity are defined as:
$\mathsf{Suf}_{d,p_{R}}({\bm{x}},{\bm{e}})=\mathbb{E}_{{\bm{x}}^{\prime}\sim p_{R}}[d(p_{\theta}(\cdot\mid{\bm{x}})\ \|\ p_{\theta}(\cdot\mid{\bm{x}}^{\prime}))]$ (5)
$\mathsf{Nec}_{d,p_{C}}({\bm{x}},{\bm{e}})=\mathbb{E}_{{\bm{x}}^{\prime}\sim p_{C}}[d(p_{\theta}(\cdot\mid{\bm{x}})\ \|\ p_{\theta}(\cdot\mid{\bm{x}}^{\prime}))],$
where $d$ is a divergence between distributions of choice, and $p_{C}$ and $p_{R}$ are interventional distributions specifying the set of allowed perturbations to the explanation and its complement, respectively. Eq. 5 are then normalised to $[0,1]$, the higher the better, via a non-linear transformation, i.e., taking $\exp(-\mathsf{Suf}_{d,p_{R}}({\bm{x}},{\bm{e}}))$ and $1-\exp(-\mathsf{Nec}_{d,p_{C}}({\bm{x}},{\bm{e}}))$. Operationally, for a given instance ($\bm{x}$}, $\bm{e}$}) sampling from $p_{C}({\bm{x}},{\bm{e}})$ equals to generating a new image where the complement of the explanation is left intact, and where perturbations are applied to the explanation. The set of allowed perturbations $p_{C}$ and $p_{R}$ can be arbitrarily defined, and different techniques are oftentimes reported to give different interpretations (Hase, Xie, and Bansal 2024; Rong et al. 2022). To avoid this confounding effect, we report the results for three different baseline perturbations, namely uniform and Gaussian noise, and black patches, along with a more advanced information-theoretic strategy named ROAD (Rong et al. 2022). Since explanations are oftentimes in the form of soft relevance scores over the entire input, a threshold is needed to tell apart relevant from irrelevant image regions. To avoid relying upon this hard-to-define hyperparameter, we aggregate the scores across multiple thresholds keeping only the best value. Therefore, for each explanation threshold value, pixels are sorted based on their relevance [^2] For sufficiency, pixels are sorted in ascending order. For necessity, in descending order. and progressively perturbed until reaching the fixed threshold value, while leaving the others unchanged. For each of those samples, we evaluate the normalized Eq. 5 where $d$ is the absolute difference in class-predicted confidence between clean and perturbed images, i.e., $|p_{\theta}(\hat{y}\mid{\bm{x}})-p_{\theta}(\hat{y}\mid{\bm{x}}^{\prime})|$, and average across the number of perturbed pixel for each threshold value. This procedure is detailed in Algorithm 1. Considering the novelty of our human driven study and the existence of such metrics, an interesting experiment is to measure if the the rating given by humans correlate with ROAD. Results, shown in Table 7 indicates a rather weak correlation. We conclude that our human scores indeed cover an aspect of explanation quality unrelated to that of perceptual quality, as predicted by Biessmann and Refiano (2021). The results cover only image-level attribution methods (see Table 5), as CBMs do not support such kinds of input-level manipulations. Our findings reinforce the idea that human evaluations and computational metrics measure complementary aspects of XAI methods. Human evaluations excel at assessing the usefulness of explanations, aligning with their primary purpose of serving a human audience. In contrast, computational metrics, such as faithfulness, focus on evaluating the alignment between the explanation and the model’s actual internal functioning. This aspect lies beyond the reach of human judgment, as humans cannot directly access or fully comprehend the internal mechanisms of the model. Evaluating the quality of an explanation typically involves estimating different and potentially orthogonal aspects of it. In addition to the perceptual quality addressed in this work, others can be numerically simulated by having access to model weights. In this additional analysis, we consider some of those aspects and measure how much they correlate with human scores.
Table 7: Pearson Correlation Coefficient (PCC) and Spearman rank Correlation Coefficient (SCC) between faithfulness computed with different methods and human scores. Values are presented with p-values in parentheses.
ROAD Black patches Uniform noise Gaussian noise
PCC SCC PCC SCC PCC SCC PCC SCC
Q1 0.029 (1e-4) -0.047 (8e-10) 0.111 (1e-47) 0.077 (1e-23) 0.037 (1e-6) -0.036 (2e-6) 0.036 (1e-6) -0.032 (3e-5)
Q2 0.030 (1e-4) -0.044 (1e-9) 0.108 (3e-45) 0.075 (2e-22) 0.034 (9e-6) -0.037 (1e-6) 0.034 (1e-5) -0.032 (3e-5)
Q3 0.035 (6e-6) 0.003 (7e-1) 0.079 (3e-25) 0.065 (2e-17) 0.029 (2e-4) 0.005 (5e-1) 0.027 (1e-4) 0.004 (6e-1)
Q4 0.033 (1e-5) -0.008 (3e-1) 0.081 (3e-26) 0.062 (1e-15) 0.030 (9e-5) -0.001 (9e-1) 0.029 (2e-4) -0.001 (9e-1)
Q5 -0.060 (8e-15) 0.003 (7e-1) -0.067 (3e-18) -0.028 (1e-4) -0.066 (6e-18) -0.025 (1e-3) -0.065 (3e-17) -0.022 (4e-3)
Q6 -0.019 (1e-2) 0.059 (1e-14) -0.053 (4e-12) -0.031 (6e-5) -0.023 (1e-3) 0.029 (1e-4) -0.022 (5e-3) 0.030 (1e-4)
Algorithm 1 Pseudo code for computing sufficiency/necessity
0: Image $\bm{x}$}, explanation $\bm{e}$}, and set of explanation-size thresholds $\mathcal{T}$.
1:values=[$\text${values}=[]
2:for each threshold $t\in\mathcal{T}$ do
3:  if computing sufficiency then
4:   Sort pixels of $\bm{x}$} in ascending order of relevance scores from $\bm{e}$}.
5:  else
6:   Sort pixels of $\bm{x}$} in descending order of relevance scores from $\bm{e}$}.
7:  end if
8:  $\text{arr}\leftarrow[]$
9:  for $i$ in range(start=1, end=$t$, step=2) do
10:   $x^{\prime}\leftarrow$ Apply the specified perturbations to the first $i\%$ sorted pixels.
11:   Append $d=\left|p_{\theta}(\hat{y}\mid\mathbf{x})-p_{\theta}(\hat{y}\mid\mathbf{x}^{\prime})\right|$ to arr
12:  end for
13:  if computing sufficiency then
14:   Append $\exp(-\text{mean}(\text{arr}))$ to values
15:  else
16:   Append $1-\exp(-\text{mean}(\text{arr}))$ to values
17:  end if
18:end for
19:Output: $\max(\text{arr})$
@@
Robustness roughly refers to how stable the explanation is to small input perturbations. Different ways to estimate it exist (Alvarez-Melis and Jaakkola 2018b; Montavon, Samek, and Müller 2018; Yeh et al. 2019; Dasgupta, Frost, and Moshkovitz 2022; Agarwal et al. 2022a). In our analysis, we focused on MaxSensitivity (Yeh et al. 2019), which applies random input perturbations to the entire image and measures the pixel-wise difference between the original explanation, and the one obtained on the perturbed sample. Formally:
$\text{MaxSensitivity}=\max\lVert{\bm{e}}-{\bm{e}}^{\prime}\lVert$ (6)
where $\bm{e}$} and $\bm{e}$$^{\prime}$ are the explanations for the original and the perturbed image, respectively. Again, different perturbation techniques can be applied, and we resort to the two simple baselines, namely Uniform and Gaussian noise. No normalization is applied, therefore the values are the higher the worse. More advanced techniques like ROAD (Rong et al. 2022) cannot be applied in this context, since the perturbation is applied uniformly over the entire image. In Table 8, we report the correlation between MaxSensitivity and human scores, outlining a non-significant correlation with the metric and some questions. Surprisingly, the most correlated questions are Q1-4, which are not requesting humans to assess the stability of the explanation, something instead partially addressed by Q5 and Q6. However, the correlation is very weak anyway, questioning any further claims.
@@
As humans have an implicit tendency to favor simple alternatives when facing a comparison between different hypotheses, providing simple and compact explanations is vital for human-machine synergy (Cowan 2001). Alternative methods for estimating the complexity of an explanation are available, from simple above-threshold counting to more advanced information-theoretic techniques (Chalasani et al. 2020; Bhatt, Weller, and Moura 2021; Nguyen and Mart´ınez 2020). To test whether those metrics are correlated to human scores, we report in Table 9 the correlation between human votes and Sparseness (Chalasani et al. 2020), which estimates explanation complexity as the Gini Index (Hurley and Rickard 2009) of the absolute values of the image attribution. The result is a metric value in the range $[0,1]$, where higher values indicate more sparseness. The computation of the Gini index is detailed in Algorithm 2.
Algorithm 2 Pseudo code for Gini coefficient calculation from Hedström et al. (2023)
0: Explanation $\bm{e}$}
[…]
6: return $\frac{\sum(2\cdot index-n-1)\cdot array}{n\cdot\sum array}$
@@
Q1 -0.137 (4e-72) -0.134 (3e-69)
Q2 -0.134 (6e-69) -0.131 (5e-66)
Q3 -0.080 (3e-25) -0.075 (9e-23)
Q4 -0.079 (6e-25) -0.073 (2e-21)
Q5 -0.085 (1e-28) -0.087 (6e-30)
Q6 0.052 (9e-12) 0.073 (9e-22)
We used the Quantus library (Hedström et al. 2023) for implementing the previous metrics, and we present the raw metric values in Table 10, aggregated by explainer and model. Overall, none of the above metrics exhibits a significant correlation with user scores.
@@
GradCAM resnet50 0.12 $\pm$ 0.20 0.92 $\pm$ 0.35 0.65 $\pm$ 0.11
GradCAM vitB 0.02 $\pm$ 0.03 1.65 $\pm$ 1.77 0.50 $\pm$ 0.26
AblationCAM vitB 0.02 $\pm$ 0.03 1.51 $\pm$ 1.19 0.74 $\pm$ 0.21
AblationCAM CLIP-zero-shot 0.02 $\pm$ 0.02 1.39 $\pm$ 0.70 0.79 $\pm$ 0.16
EigenCAM resnet50 0.12 $\pm$ 0.21 0.93 $\pm$ 0.47 0.79 $\pm$ 0.07
EigenCAM vitB 0.02 $\pm$ 0.03 1.03 $\pm$ 0.33 0.58 $\pm$ 0.06
EigenCAM CLIP-zero-shot 0.02 $\pm$ 0.02 0.78 $\pm$ 0.38 0.56 $\pm$ 0.09
EigenGradCAM resnet50 0.12 $\pm$ 0.20 0.96 $\pm$ 0.57 0.78 $\pm$ 0.07
EigenGradCAM vitB 0.02 $\pm$ 0.03 2.40 $\pm$ 1.74 0.83 $\pm$ 0.19
EigenGradCAM CLIP-zero-shot 0.02 $\pm$ 0.03 1.26 $\pm$ 0.61 0.72 $\pm$ 0.14
FullGrad resnet50 0.11 $\pm$ 0.19 0.48 $\pm$ 0.15 0.40 $\pm$ 0.06
FullGrad vitB 0.02 $\pm$ 0.02 1.08 $\pm$ 0.48 0.38 $\pm$ 0.06
GradCAM CLIP-zero-shot 0.02 $\pm$ 0.02 1.21 $\pm$ 0.66 0.57 $\pm$ 0.17
GradCAMPlusPlus resnet50 0.12 $\pm$ 0.20 0.71 $\pm$ 0.26 0.60 $\pm$ 0.10
GradCAMPlusPlus vitB 0.02 $\pm$ 0.02 2.33 $\pm$ 2.17 0.59 $\pm$ 0.28
GradCAMPlusPlus CLIP-zero-shot 0.02 $\pm$ 0.02 1.46 $\pm$ 0.93 0.63 $\pm$ 0.24
GradCAMElementWise resnet50 0.11 $\pm$ 0.20 0.72 $\pm$ 0.25 0.58 $\pm$ 0.10
GradCAMElementWise vitB 0.02 $\pm$ 0.02 1.20 $\pm$ 0.40 0.59 $\pm$ 0.09
GradCAMElementWise CLIP-zero-shot 0.02 $\pm$ 0.02 0.71 $\pm$ 0.28 0.38 $\pm$ 0.08
HiResCAM resnet50 0.12 $\pm$ 0.20 0.91 $\pm$ 0.35 0.65 $\pm$ 0.11
HiResCAM vitB 0.02 $\pm$ 0.03 1.92 $\pm$ 1.13 0.71 $\pm$ 0.24
HiResCAM CLIP-zero-shot 0.02 $\pm$ 0.02 1.43 $\pm$ 0.49 0.69 $\pm$ 0.13
LIME resnet50 0.15 $\pm$ 0.22 0.49 $\pm$ 0.06 0.12 $\pm$ 0.04
ScoreCAM resnet50 0.12 $\pm$ 0.21 0.78 $\pm$ 0.31 0.59 $\pm$ 0.10
ScoreCAM vitB 0.02 $\pm$ 0.03 1.22 $\pm$ 0.65 0.53 $\pm$ 0.18
XGradCAM resnet50 0.12 $\pm$ 0.20 0.88 $\pm$ 0.34 0.65 $\pm$ 0.11
XGradCAM vitB 0.02 $\pm$ 0.03 1.37 $\pm$ 0.30 0.62 $\pm$ 0.07
XGradCAM CLIP-zero-shot 0.02 $\pm$ 0.02 1.30 $\pm$ 0.20 0.59 $\pm$ 0.06
DeepFeatureFactorization resnet50 0.11 $\pm$ 0.18 0.95 $\pm$ 0.52 0.34 $\pm$ 0.12
DeepFeatureFactorization vitB 0.02 $\pm$ 0.03 0.63 $\pm$ 0.23 0.28 $\pm$ 0.05
DeepFeatureFactorization CLIP-zero-shot 0.02 $\pm$ 0.03 0.65 $\pm$ 0.31 0.33 $\pm$ 0.08
@@
BCos resnet50-bcos 0.15 $\pm$ 0.24 0.88 $\pm$ 0.31 0.50 $\pm$ 0.09
LIME CLIP-zero-shot 0.02 $\pm$ 0.03 0.58 $\pm$ 0.09 0.16 $\pm$ 0.05
SHAP resnet50 0.13 $\pm$ 0.21 1.37 $\pm$ 0.46 0.50 $\pm$ 0.10
SHAP vitB 0.02 $\pm$ 0.03 1.05 $\pm$ 0.45 0.40 $\pm$ 0.09
SHAP CLIP-zero-shot 0.02 $\pm$ 0.03 1.19 $\pm$ 0.38 0.46 $\pm$ 0.10
AblationCAM resnet50 0.12 $\pm$ 0.19 1.44 $\pm$ 0.48 0.71 $\pm$ 0.11
B.3 Dataset analysis